### v5

Abstract
Evaluating early design concepts is crucial as it impacts quality and cost. This process is often hindered by vague and uncertain design information. This article introduces the SysML-based Simulated-Physical Systems Modeling Language (SPSysML). It is a Domain-Specification Language for evaluating component reusability in Cyber-Physical Systems incorporating Digital Twins and other simulated parts. The proposed factors assess the design quantitatively. SPSysML uses a requirement-based system structuring method to couple simulated and physical parts with requirements. SPSysML enables DTs to perceive exogenous actions in the simulated world.
SPSysML validation is survey- and application-based. First, a robotic system for an assisted living project was developed. The integrity of simulated and physical parts of the system is improved using SPSysML-based quantitative evaluation. Thus, more system components are shared between the simulated and physical setups. The system was deployed on the physical robot and two simulators based on the Robot Operating System (ROS) or ROS2. SPSysML was used by a third-party developer and was assessed by him and other practitioners in a survey.
keywords:
[…]
1 Introduction
Simulation is widely used in state-of-the-art cyber-physical systems (CPS) development procedures. Recent papers refer to Model-in-the-loop (MIL) [1], Software-in-the-loop (SIL) [2], Hardware-in-the-loop (HIL) [3] and Rapid Control Prototyping (RCP) [4] techniques. They are used selectively or are composed in sequence, e.g. verification and validation steps in a systems development approach called V-model [5]. Each of the techniques requires a simulation of the system parts. Some systems utilise the Digital Twin (DT) [6, 7] concept. They employ an accurate simulation of a system part for, e.g. Swapping malfunctioned parts, Energy consumption analysis, Technology integration, or Real-time monitoring.
The system’s parts (software and hardware) interact with the real world, while others interact with the simulated world. We define the set of the parts used in the real world as the physical embodiment and the set of the parts used in the simulated world as the simulated embodiment. The DT is a part of the simulated embodiment mirroring a part of the physical embodiment. DT concept is used in numerous domains and applications [8]. They are systematically analysed by the authors of [9].
Robots comprise multiple devices and their controllers. In particular, autonomous mobile robots require a complex navigation system featuring multiple closed control loops, e.g. drive controllers, trajectory controllers, and Simultaneous Localisation and Mapping (SLAM). Furthermore, in a complex mobile manipulator system [10], the navigation system is integrated with manipulation control to execute user requests (like object transportation). In some cases, only a fraction of such a complex system must have DTs. In other cases, if the robot system only has the simulated embodiment, it is a demonstrator of a future product.
To clear up the taxonomy of systems featuring simulated and physical parts, we introduce the Simulated-Physical System (SPSys) concept (Fig. 1). This kind of system consists of at least one physical/simulated embodiment and a shared controller. If it has both embodiments, the simulated part can be a DT of a physical part. If it has only the physical embodiment, it is a CPS; if it has only the simulated embodiment, it is a simulator.
Figure 1: The idea of Simulated-Physical System
Developing reliable systems, especially complex ones like robot systems, requires comprehensive unit and integration testing. Some CPS parts can be tested with simulated hardware only; therefore, additional parts that do not comprise the operational system are required (e.g. human simulator in a social robot case). Comprehensive testing is complex in test case specification and time-consuming in test implementation and execution. From this perspective, software reusability is a key to fast development of complex, reliable systems.
We aim to:
  • Enhance system requirements traceability for simulated and physical system parts thanks to the introduction of profiled-based requirements of Simulated-Physical Systems, including multiple Digital Twins for procedural system structuring based on the requirement types,
  • Evaluate the level of component re-use between DT and its Physical Twin (PT) and between different system setups, improving the system’s reliability and resulting in more accessible and faster testing,
  • Enable DTs to observe exogenous actions in simulation. The actions can be executed by any agent outside the physical system, e.g. a human moving objects in a service robot environment. Such an action must be implemented in the simulated world to make DT operate in the exact environment as its physical counterpart.
To reach these goals, we propose a Domain-Specific Language (DSL) named Simulated-Physical System Modeling Language (SPSysML). SPSysML models the requirements and the SPSys, in particular, SPSys consisting of simulated parts (used either in the system development or in the operational setup as a DT/simulator). Based on SPSysML, we propose a requirement-based SPSys structuring method for enhancing requirement readability in simulated and physical system parts. To measure the inter-embodiment integrity of the system, we propose quantitative design evaluation factors for SPSys designs. The factors are used during system development to maximise the shared controller, minimise the variety of the system parts and improve simulated-physical integrity. Our goal originates from the conclusion of [11]: The digital models are mainly used to examine product performance(…). However, how to optimise the use of those models to enhance the design process and design collaboration still needs to be investigated.
In this article, we describe related work concerning mixing simulated-physical setups and meta-models for CPS (Sec. 2) and state our work’s novelty and its result regarding the related works (Sec. 3). Subsequently, we define SPSys requirements model in Sec. 4.1, the meta-model in Sec. 4.2. In Sec. 5, we analyse the proposed evaluation factors of SPSys design and describe characteristics of a system scored edge case values of the factors. Finally, we validate our methodology in complex SPSys development for the INCARE project (Sec. 6.1) and by robotics practitioners assessment 6.2.
(a) Transport task
(b) Fall assistance task
Figure 2: TIAGo robot in INCARE tasks execution
Table 1: Related works comparison with this work
System development method EIV Formal meta-model SCP SE System structure configurability DE E QE MDT Purpose
Digital Mockup [12] - - ES - Single structure - - - Real time (RT) simulation for HIL testing
C2PS [13] - Own meta-model Formal and detailed BS Single adaptive - - N/D Digital Twin in the cloud
[14] Dolev-Yao [15] Formal and detailed - Single structure - - Cyber-security, physical-digital twin synchronisation
DEVSRT [16] - DEVS [17] General - Single structure - - - - Simulation to embedded continuous development
aRD[18] - - General - Multiple setups - - - RT & NonRT parts integration
HMLF [19] - - ES - Single structure - - - - Simulation-driven ML
RSHPN [20] - RSHPN /Petri net ES DI Single structure - - - Control deadlock check, code generation
SPSysML SPSysML /SysML ES / Formal and detailed SPE Multiple setups Structure evaluation
EIV – Embodiment integrity verification, SCP – System creation procedure, SE – Structure evaluation, DE – supports simulated and physical components, DI – Deadlock identification, E – If uncontrolled agents of the physical environment (link humans) can be modelled in the simulation, QE – If quantitative evaluation factors for system design are proposed, MDT – If a multi Digital Twin system model is presented, ES – Example system as the method application procedure, BS – The best selection from the previously designed, SPE – Simulated-Physical design evaluation, N/D – Possible, however, not defined,
2 Related work
2.1 Literature review constraints
The key related works are described in Table 1, where their crucial features, including goals, are compared with SPSysML. Other works are briefly described in the following sections.
2.2 Simulated-physical system development
In [18], the authors note the robot development process features parallel tasks conducted by heterogeneous R&D teams. The authors suggest system simulation enables rapid prototyping and an iterative development process. They propose the agile Robot Development (aRD) concept facilitating the integration of the hard real-time part and the non-real-time part of the system utilising a Matlab/Simulink toolchain. The sim2real problem is known and mentioned in the recent research results, e.g. in the domain of self-driving cars [21]. However, to our knowledge, none of the existing works specifies a robotic system meta-model with integrity evaluation for mixed simulated-physical setups of CPS.
There are reviews on CPS [22] and DT [11] development. The CPS-related works focus on controlling a physical object and its model in a simulation environment with partially or fully simulated hardware. For example, the authors of [14] use the DT approach to run the digital embodiment in a safe virtual machine and confront the physical and digital embodiments to spot anomalies caused by a cyber-attack. Another work [19] employs simulation-driven machine learning for robots. However, none of the works proposes a system structure allowing exogenous actions execution in the simulated world.
2.3 System design evaluation
Early-stage product development challenges systems engineers in designing accurate structures and architectures. Evaluating design concepts is crucial as it mightily influence both the quality and cost of the final product.
Tools and models supporting designers in multiattribute utility analysis are the core of this research area. An example Methodology for the Evaluation of Design Alternatives (MEDA) [23] was presented already in 1991. It is based on the subjective designer’s view of the design alternative attributes, e.g. utility. Nowadays, customers play a crucial role in the design process. The authors of [24] tackle the problems that arise in this situation. Priorities are set, and the best design alternative(s) are chosen using the proposed comparison rules. Multi-criteria decision-making is resolved as a constrained multi-attribute optimisation solution [25] or by integrating rough sets in handling vagueness with grey relation analysis [26]. The system’s integrity investigated in our work is another system attribute; thus, our evaluation factors values can provide unbiased, objective input to the design evaluation methods and make the design decision aware of the component reuse level between the system’s simulated and physical embodiments.
2.4 Meta-models for CPS
Engineers use a specific language to plan, conceptualise, and specify the system. Currently, the state-of-the-art approach is a model-based approach utilising a DSL. There are DSLs supporting verification and testing of, e.g. industry 4.0 plants [27], or agent-based computational systems [28]). Unified Modeling Language (UML) is the most known language; however, System Modeling Language (SysML) [29] is proposed by the Object Management Group (OMG) to support the design, analysis and verification/validation of complex systems comprising software and hardware components. Different Model Driven Engineering approaches exist in the robotics domain [30]. For instance, Embodied Agent-Based cybeR-physical control systems modelling Language (EARL) [31] is SysML specialisation for robotics which allows analysis and specification of the robotic system properties. It is based on the Embodied Agent approach [32]. There are various design analysis tools based on this approach: top-down communication-based system decomposition [33], and Petri net-based design [34] for control deadlocks analysis [20]. It was applied to robotic and multi-modal human-machine interface [35] systems. This article extends the Embodied Agent approach to the simulated and DT-related components domain. We manage this extra complexity with the proposed requirement-based design method and quantitative system design evaluation.
3 Contribution and result validation
None of the models presented in Sec. 2 specifies SPSys in general, especially including a physical part without equivalent Digital Twin, simulated parts without equivalent PT, or hybrid SPSys including Twins and non-Twin components. The key novel features delivered by SPSysML are:
  • Integrity evaluation with quantitative factors based on SysML-based model,
  • Guidance for the simulated-physical integrity increase measured by the quantitative factors,
  • Requirement-based system composition method for straightforward requirement satisfaction tracing,
  • Specification of Simulated-Physical System including multiple physical and simulated parts and multiple Digital Twins,
  • Reflection of the dynamic environment in simulation to enable Digital Twins to observe exogenous actions,
  • A survey-based assessment of Domain-specification Language,
  • Validation of a system design evaluation method in robotics by system implementation in ROS and ROS2.
Based on our experience in SPSys development (TIAGo robot [36, 37, 38], Velma robot [39], IRP6 robot [^1] Real: https://youtu.be/wJpFcy99Gh0, Simulation: https://youtu.be/BjwcbSdouHw [40, 41]) and the literature analysis, we propose the SPSysML allowing software integrity evaluation between the system embodiments, and further by iterative system design, maximise the between embodiments integrity. Complex system structuring is a broad topic, and solutions are utilised using various methods and strategies, e.g., top-down, middle-out, and bottom-up. In [33], we describe a binary communication-focused top-down approach for robotic systems. This article proposes a requirement-based bottom-up system structuring method customised for SPSys.
SPSysML comprises a componet-based Platform Independent Model (PIM), thus, can be applied to any SPSys. As the system is developed, e.g. with Robot Operating Framework (ROS), it becomes a Platform Specific Model (PSM) that can be launched in the specified setups. PSM implementation utilises platform-specific tools and software libraries. Therefore, to validate our approach, we design, implement, test and deploy a specific SPSys. It uses the TIAGo service robot [42] for the INCARE (Integrated Solution for Innovative Elderly Care) project [^2] http://aal-incare.eu/, robot application demonstration (simulated/real) can be found accessing Demo / RAPP Store Demo. We integrate extended voice interface and additional devices to serve the elderly, e.g. in object transportation (Fig. 2(a)) and fall assistance (Fig. 2(b)). The system implementation is component-based and uses frameworks like ROS [43, 44], ROS2 [45], OROCOS [46], that are a standard open-source robot control frameworks. We use two open-source simulators Gazebo [47] (in particular Gazebo_ros_control package [48]) and O3DE [49] for DT implementation. Besides validation by application, we present results of SPSysML assessment among systems engineering practitioners. To make the SPSys development easier for the community, we share the SysML profiles, the SPSysML meta-model and the example INCARE system model [^3] https://github.com/RCPRG-ros-pkg/spsysml. All diagrams describing SPSysML and the INCARE system are also shared [50].
4 Simulated-Physical System
Simulated-Physical Systems Modeling Language (SPSysML) is based on SysML and EARL. In the following sections, we describe SPSysML, which defines a stereotype-based meta-model of requirements and system structure. Description of a DSL requires a notation for multiple blocks or instances of a stereotype (e.g. two instances of <<Agent>>). For this purpose, we append the stereotype with s (e.g. <<Agent>>s). We refer to an instance of a stereotype using the part-name <<stereotype>> symbol.
4.1 SPSysML – requirements meta-model
We define a stereotype-based model of requirements (Fig. 3). The stereotypes explicate SysML requirement diagrams used for designing SPSys with SPSysML.
Figure 3: Model of system requirements defined for SPSys
SPSysML specifies the requirements model as SysML profile. The requirements can be defined based on various premises. In particular, they can derive from user requirements [51]. The SPSysML profile defines structural, functional, configurational and environment requirements. Other requirement types can be added; however, these are used in system setups’ specification and requirement-based system composition. Auxiliary sequence diagrams presenting the concept of the system behaviour and use case diagrams are useful for defining requirements. Before SPSys requirements specification, its environment must be analysed. In SPSysML, we distinguish <<ExogAgentReq>> being a part of the environment requirements that specify exogenous agents interacting with the world alongside the system. The system functions given by <<FunctionalReq>> stereotyped requirements are classified into:
  • simulated-physical, which require perceiving or affecting the simulated/physical world— <<SimPhyFunReq>>,
  • computational, which do not interact directly with any simulated/physical world— <<CompFunReq>>.
Each <<FunctionalReq>> requirement must be performed by a system part; thus <<SysPartReq>> stereotyped requirements must be specified. As shown in Fig. 3, they must satisfy, verify, and derive from their origin— <<FunctionalReq>>. In SPSysML, the system structure is based on the requirements of <<SysPartReq>>s stereotype. Each <<SysPartReq>> is classified as elementary <<HardwareReq>> or one of more general <<PhyPartReq>>, <<SimPartReq>> or <<SimPhyPartReq>>.
  • <<PhyPartReq>> determines a part interacting only with the Physical World (even during the system development, e.g. during parallel development of the connected parts), and its simulated embodiment is not required (or cannot be created) during the system development.
  • <<SimPartReq>> specifies a part interacting only with the Simulated World (e.g. mock-ups or demonstrators). <<SimPhyPartReq>> interacts with both Worlds (e.g. realised with a pair of DT and PT).
  • <<HardwareReq>>s are in a satisfy relationship with <<FunctionalReq>>s. This means hardware requirements specified with a <<HardwareReq>> enable the realisation of the functionality specified with the given <<FunctionalReq>>.
The configurational stereotypes specify if a system part (<<SysPartReq>>) or a system function (<<FunctionalReq>>) is obligatory to launch the system in all of its setups. Configurational stereotypes are orthogonal to structural and functional stereotypes; thus, functional and structural requirements may be either <<ObligatoryReq>> or <<OptionalReq>>. However, a <<SysPartReq>> realising a <<FunctionalReq>> derives its configurational stereotype from the functional requirement. Each <<OptionalReq,FunctionalReq>> gives the system more setups, and the number of the setups equals $2^{n}$, where $n$ is the number of <<OptionalReq,FunctionalReq>>. As a result,
  • <<ObligatoryReq,FunctionalReq>> requirements define the core system’s functions,
  • <<OptionalReq,FunctionalReq>> requirements manage the system’s versatility by dividing the system into system setups with different functionality,
  • <<ObligatoryReq,SysPartReq>> requirements define the system’s core components,
  • <<OptionalReq,SysPartReq>> requirements define the system’s components for different functionality setups.
4.2 SPSysML – structure of SPSys
SPSysML derives from EARL [31] version 1.3 [52]. In a SPSys we differentiate three specialisations of the <<Agent>> stereotype defined in EARL (Fig. 4):
  • Physical Agent (<<PhyAgent>>) – Runs only in the physical embodiment (in particular Agent interacting with or sensing real world),
  • Simulation Agent (<<SimAgent>>) – Runs only in the simulated embodiment (in particular Agent interacting with or sensing simulated world),
  • Hybrid Agent (<<HybAgent>>) – Runs in both embodiments and computes without interacting or perceiving any world.
Figure 4: Simulated-Physical System composition
The Groups of Agents called World Mirror Group of Agents aggregate <<SimAgent>>s that are Digital Twins of the Agents in the Physical World that are not controlled by the physical embodiment of the system. They execute exogenous actions in the Simulated World. Groups of this type exist if the system works in a dynamic environment (e.g. an environment with human inhabitants). For example, a World Mirror Group of Agents modifies the Simulated World as humans do in the Physical World.
An agent is composed of Subsystems of different types as specified in EARL. Types of Subsystems are defined in 4.2.2. We use a Group of Subsystems (<<GpSubsys>>) to gather Subsystems with a specific common properties, and a Group of Agents (<<GpAgents>>) to organise the Agents (composed of Subsystems) cooperating for a defined aim in the system.
4.2.1 Digital Twins in SPSys
Integrity between SPSys embodiments is crucial. Therefore, creating DT of a physical parts is advisable and common. We define the mirror relationship between <<PhyAgent>> and <<SimAgent>> to model the relationship between DT and its PT. To enable multi-agent DT for a single <<PhyAgent>> and vice versa, we introduce <<MirrPhyGpAgents>> and <<MirrSimGpAgents>>. They aggregate Agents of different stereotypes (Fig. 5).
Figure 5: SPSys composed of 0…* Digital Twins (+dt) and Physical Twins (+pt) realised with <<MirrSimGpAgents>> and <<MirrPhyGpAgents>> accordingly
The definition of the mirror relationship is as follows:
Definition 4.1 (Mirror relationship)
Two Groups of Agents are said to be in a mirror relationship if their input buffers and goals are the same and affect Simulated and Real Worlds congruently. The relationship is an association between Digital and Physical Twins.
It is worth noting that the same stimuli of the mirroring Groups cause corresponding reactions in specific worlds, in the Simulated World for Mirror Simulation Group of Agents and in the Physical World for Mirror Physical Group of Agents. The particular reaction results from the requirements of the specific system and does not need to be identical between the Agent Groups that mirror each other. A stimuli for a Group of Agents is the input from other Agents or the environment through the interfaces. Agent’s reaction to a stimuli manifests as the Agent state change (including changes in memory or output data).
Example DT-PT pair
Considering a pan-tilt head as an example, <<MirrSimGpAgents>> is the simulated head, and <<MirrPhyGpAgents>> is the real one. The stimuli of this DT-PT pair include the light perception from the according world (simulated/physical) and the control for the head joints. The common reaction is data in their output interface. The DT (<<MirrSimGpAgents>>) for a pan-tilt head development is a convenient tool, e.g. for servovision implementation or joint constraints evaluation. SPSys may consist of multiple such pairs, e.g. manipulator and mobile base in a service robot case.
Constaints and guides for DT-PT pair
The mirror relationship does not propagate automatically to all the Agents aggregated in the mirroring Groups, so agents from <<MirrSimGpAgents>> do not mirror all agents of mirroring <<MirrPhyGpAgents>> and vice versa.
The number of Agents outside DT/PT pairs should be minimised to maximise coverage of simulation-based testing and analysis. Additionally, the complexity and quantity of mirroring Agent Subsystems should be minimised to boost the modularity and integrity of the system. Otherwise, the system could be composed of just two mirroring agents. If a computational functionality is required in both embodiments, a <<HybAgent>> should be designed for this purpose.
4.2.2 Subsystem types
The Agents are built with Subsystems of different types. In EARL, the central computational part of an Agent is the Control Subsystem <<ContSubsys>>. An Agent communicates with other Agents using communication buffers of its <<ContSubsys>>. To percept the environment, an Agent uses Real Receptors (<<RealRec>>s), and to aggregate and preprocess stimuli, it uses Virtual Receptors (<<VirtRec>>s). To affect the environment, an Agent uses Real Effectors (<<RealEff>>s). To preprocess <<ContSubsys>> commands to signals for <<RealEff>>s, it uses Virtual Effectors (<<VirtEff>>s). Detailed description and SysML diagrams of Agent structure and behaviour are published in [31].
SPSys interacts with both the simulated and physical world; thus, SPSysML differentiates between the types of the above Subsystems— Simulated, Physical, and Simulated-Physical. To achieve maximum integrity between the simulated and physical embodiments, all <<ContSubsys>>s should be Simulated-Physical (<<SimPhyContSubsys>> stereotype) and constitute the general concept of the shared controller, recall Fig. 1. For iterating design purposes, embodiment-specific Control Subsystems concepts (<<PhyContSubsys>> and <<SimContSubsys>>) may be helpful. Virtual and real effectors/receptors can be Simulated or Physical. In SPSysML we group them in four <<GpSubsys>> that aggregate receptors and effectors for each embodiment— simulated and physical (Fig. 6, Fig. 7). Real effectors/receptors are hardware parts, and Virtual effectors/receptors are their drivers.
In complex systems, one <<SimPhyContSubsys>> may communicate with many Simulated/Physical Virtual Receptors and Effectors. In this case, the <<SimPhyContSubsys>> is reused in <<SimAgent>> and <<PhyAgent>>, and both utilise embodiment-specific hardware and drivers. The data flow and communication links for <<SimPhyContSubsys>> interacting with Simulated and Physical Worlds are shown in (Fig. 8).
Figure 6: Embodiments of Real Effectors and Receptors
Figure 7: Embodiments of Virtual Effectors and Receptors
Figure 8: Interfaces between Subsystem Groups on example, where two agents share <<SimPhyContSubsys>>
Agent composition constraints
Each Agent type defined in SPSysML aggregates a particular number of Subsystems of a specific type (Fig. 9).
(a) Physical Agent
(b) Simulation Agent
(c) Hybrid Agent
Figure 9: Subsystems aggregated by Agent classes
It should be noted that the Basic EARL meta-model defines only one <<ContSubsys>> in an Agent, and inter-agent communication is handled only by the <<ContSubsys>>. In SPSysML, we differentiate between specific types of <<ContSubsys>> for the embodiments; still, an Agent can aggregate just one specialisation of <<ContSubsys>>.
System setups composition
SPSys can be deployed in various setups (e.g. testing setups). The setups are a Group of Agents fulfilling the system’s functionality in the setup. For the system setup definition, we use <<SysSetupGpAgents>> to specify the Group of Agents working in the setup. The set of agents composing <<SysSetupGpAgents>> is derived from the system’s requirements and test scenarios.
Example Physical embodiment realisation
Subsystems in a <<PhyHardGpSubsys>> (sensors/actuators) communicate with their drivers via various interfaces (e.g. Linux kernel driver for Inter-Integrated Circuit (I2C) or network interface). <<PhyDriverGpSubsys>>s take data from these interfaces, aggregate it and expose their interface for a <<ContSubsys>> executing the logic of the Agent.
Example Simulated embodiment realisation
Subsystems in <<SimHardGpSubsys>> expose Application Programming Interface (API) classes, defined in the simulation environment (e.g. gazebo::ModelPlugin for <<SimRealEff>> and gazebo::SensorPlugin for <<SimRealRec>> in Gazebo444a popular simulation environment used in the DARPA Robotics Challenge, in July of 2013). Subsystems of <<SimDriverGpSubsys>> connect to these interfaces using the API, aggregate sensor data, and respond <<ContSubsys>>s control.
The connections between the groups are shown in Fig. 8.
5 The design evaluation factors
We recognise the following integrity factors classified for design analysis in two scopes:
  • System-wide:
    • Controller integrity factor (IIF=$\frac{c^{dcs}}{c^{All}}$), where $c^{dcs}$ and $c^{All}$ are the cardinalities of <<SimPhyContSubsys>>s and all system <<ContSubsys>>s accordingly,
    • Driver generalisation factor (DGF=$\frac{r_{u}}{r}$), where $r_{u}$ is the count of Real Subsystems aggregated in an Agent controlled by a <<SimPhyContSubsys>> and $r$ is the count of all Real Subsystems in the system,
    • Digital Twin coverage (DTC=$\frac{a_{P}^{m}}{a_{P}^{All}}$), where $a_{P}^{m}$ is the count of <<PhyAgent>> aggregated in a <<MirrPhyGpAgents>> being a PT of a DT (<<MirrSimGpAgents>>), and $a^{P}_{All}$ is the count of all <<PhyAgent>> in the system,
  • DT/PT pair-wide:
    • Mirror integrity factor (MIFn=$\frac{c_{n}^{dcs}}{c_{n}^{All}}$), where n is the considered pair of mirroring <<MirrPhyGpAgents>> and <<MirrSimGpAgents>> composing one DT/PT pair, $c_{n}^{dcs}$ and $c_{n}^{All}$ are the counts of <<SimPhyContSubsys>>s and all <<ContSubsys>>s accordingly in nth DT.
Below, we present the interpretation of the evaluation factors and the correlation between the factor values and the SPSys features. We do it by describing the characteristics of edge case SPSys that scores maximum or minimum values of the design evaluation factors:
  • IIF – is a share of the software controller common between the embodiments. Virtual Subsystems are not considered, as their count may be related to the Real Subsystem counts in each embodiment. It is maximised by the reduction of the number of <<SimContSubsys>>s and <<PhyContSubsys>>s in favour of an inter-embodiment <<SimPhyContSubsys>>s, The higher IIF is, the more software components are shared between the simulated and physical embodiments. At maximum (IIF=1), all hardware abstract parts of the system are common.
    • IIF = 0: There are no <<SimPhyContSubsys>>, only <<PhyContSubsys>> or <<SimContSubsys>>. The system’s parts in simulated and physical embodiments are disjunctive. Simulation-based testing is not possible. The system’s functions in the simulation may be completely different from those in the physical embodiment.
    • IIF = 1: All Control Subsystems are <<SimPhyContSubsys>>, and there are no <<PhyContSubsys>> or <<SimContSubsys>>. This means all hardware abstract parts of the system are common between its embodiments, and the coverage of simulation-based testing is maximised and allows integration testing in simulation.
  • MIFn – is a share of the system parts common between Physical and Digital Twins composing the nth twin pair (managing given <<SimPhyPartReq>>). It is similar to IIF but within the scope of $n^{th}$ pair of Physical and Digital Twin. Tips for maximisation of the MIFn factor are:
    • extraction of common functions as <<HybAgent>>s from <<MirrPhyGpAgents>> and <<MirrSimGpAgents>>
    • and/or redesign of interfaces between a <<SimContSubsys>> and <<SimDriverGpSubsys>> and between a <<PhyContSubsys>> and <<PhyDriverGpSubsys>> to emerge a common <<SimPhyContSubsys>> from the <<SimContSubsys>> and <<PhyContSubsys>>,
  • DGF – is a share of Real Subsystems (hardware) controlled by <<SimPhyContSubsys>>s (shared controller). It expresses hardware control integrity between the embodiments.
    • DGF = 0: All Hardware parts are controlled by embodiment-specific Control Subsystems. The causes of this depend on a specific case:
        For Physical Hardware without a DT, it means the interface to hardware is embodiment-specific; thus, extending the system with a DT of the hardware is complicated and would require adding simulation-specific <<SimContSubsys>>.
        For Physical Hardware with a DT, it means the Hardware Drivers interface of Physical and Digital Twins differ, and the software using the interface differs between the embodiments. This means the system part designed as DT of the Physical hardware is not a proper DT.
    • DGF = 1: All Hardware parts are controlled by embodiment-abstract Control Subsystems; thus, the Agents managing hardware are interchangeable between the embodiments, or future Digital/Physical Twin integration for Physical/Simulated Hardware is straightforward.
  • DTC – is a share of hardware and its controllers mirrored with a DT. Its increase boosts coverage of simulation-based testing of hardware controllers and system robustness utilising the DT concept. If DTC = 0, there are no DTs in the system; if DTC = 1, all Physical Hardware parts have DTs.
Based on the factors’ values, one can evaluate the system design in terms of:
  • Safety and hardware independence during software testing – based on IIF (for system scope), MIFn (for $n^{th}$ DT),
  • Simulation-based testing and failure examination/prediction of the system parts – based on DTC
  • Inter-embodiment integrity of hardware controllers and readiness for simulation-based hardware testing – based on DGF.
Maximisation of these factors is not always required, and the optimisation goal can be set at a different point in the factors’ space. The goal depends on the specific system requirements. However, the factors’ values inform the designer about the inter-embodiment integrity of the system design, so her/his decision is conscious.
6 SPSysML Validation
First, we describe the SPSysML validation by application. It includes the requirement-based system composition, design evaluation, and applicability in the system development procedure 6.1. Next, we share the results of the SPSysML assessment done among systems engineering practitioners, including a third-party SPSysML user 6.2.
6.1 Example system design and analysis
SPSysML was used in developing complex SPSys utilising a service robot for the INCARE project. This is the primary validation of SPSysML. The INCARE system idea and requirements are published in [53]. The framework model developed to manage robot tasks is published in [54, 55]. Developing a complex system requires different implementations of its parts in various development phases. In INCARE, we use some commercial products like the TIAGo robot with its control system. We develop and integrate new parts into the system (e.g. human fall detector and TIAGo audio interface extension). We use and modify the TIAGo robot in this project; thus, we specify its hardware and controller as a part of the INCARE system specification. We present the result of the SPSysML application in the following SPSys development tasks.
Requirement engineering
In these steps, we use the requirement model defined in SPSysML. We specify the structural and functional requirements based on the general requirements of the INCARE project. The general and the robot-specific requirement diagrams are shown in Fig. 10.
(a) INCARE structural requirements
(b) The Robot requirements, where TTS and STT are text-to-speech and speech-to-text functions
Figure 10: Example part of the INCARE requirements
[…]
Exogeneity identification
The robot in the INCARE project coexists with humans being exogenous agents from its perspective; thus, we utilise WorldSync <<WorldMirrGpAgents>> to manage humans in the Simulated World. Detailed specification and realisation of WorldSync <<WorldMirrGpAgents>> is available in [56].
System composition
System structuring resulted in the system being composed of 5 embodiment-specific Agents, one for each system part:
  • TIAGo <<SimAgent>>, TIAGo <<PhyAgent>>,
  • FallDetector <<SimAgent>>, FallDetector <<PhyAgent>>,
  • SmartHome <<PhyAgent>>.
Additionally, there are 3 Hybrid Agents, one for each computational function of the system:
  • ComplexTaskExecution <<HybAgent>>,
  • Talker <<HybAgent>>,
  • FakeAudio <<HybAgent>>.
There are two DT-PT pairs:
  • one resulting from the requirement Robot <<SimPhyPartReq>>:
    • Robot <<MirrSimGpAgents>> being a DT composed of TIAGo <<SimAgent>>, FakeAudio <<HybAgent>> and Talker <<HybAgent>>,
    • Robot <<MirrPhyGpAgents>> being PT composed of TIAGo <<PhyAgent>> and Talker <<HybAgent>>.
  • one resulting from the requirement Fall Detector <<OptionalReq,SimPhyPartReq>>, FallDetector <<SimAgent>> as DT and FallDetector <<PhyAgent>> as PT.
Agent decomposition
To provide an example, we describe the final decomposition of two TIAGo <<SimAgent>> realisations (O3deTIAGo (Fig. 11) being the robot simulator implemented in the O3DE simulator and GazeboTIAGo (Fig. 12) implemented in Gazebo) and TIAGo <<PhyAgent>> (Fig. 13) which mirrors the previous ones. Each robot hardware component is specified as either <<PhyRealRec>>, <<PhyRealEff>>, <<SimRealRec>>, or <<SimRealEff>>. O3deTIAGo: TIAGo <<SimAgent>> integrates the system with O3DE simulation environment; thus, <<SimRealRec>>s and <<SimRealEff>>s are core O3DE components interacting with the simulated world, called gems. GazeboTIAGo: TIAGo <<SimAgent>> integrates the system with Gazebo simulation environment; thus, <<SimRealRec>>s and <<SimRealEff>>s expose gazebo::SensorPlugin and gazebo::ModelPlugin interfaces accordingly. As the physical robot, we use PAL Robotics’ TIAGo; thus, <<PhyRealRec>>s, <<PhyRealEff>>s interfaces are adequate Linux Kernel drivers managing communication with the devices. In GazeboTIAGo: TIAGo <<SimAgent>> case, the Simulated Drivers connect to the gazebo::SensorPlugin and gazebo::ModelPlugin interfaces and expose the robot state information and typical ROS topics/services (e.g. JointStateInterface and EffortJointInterface for MobileBaseController <<SimVirtEff>> [^5] This ¡¡VirtEff¿¿ is based on gazebo_ros_control package: https://classic.gazebosim.org/tutorials?tut=ros_control and /scan ROS topic for lidar <<SimVirtRec>>). If Simulated World is the Gazebo environment, <<SimVirtEff>>s and <<SimVirtRec>>s are usually implemented as Gazebo Plugins. <<PhyDriverGpSubsys>> connects to Linux Kernel drivers and, as a whole Group, exposes to <<ContSubsys>> identical interfaces as <<SimDriverGpSubsys>>. The diagrams of the TIAGo robot agents (simulated and physical) show their common RobotIf <<SimPhyContSubsys>>; however, the physical robot and Gazebo simulator run ROS1: RobotIf <<SimPhyContSubsys>>, and O3DE simulator runs ROS2: RobotIf <<SimPhyContSubsys>>.
Figure 11: IBD of O3deTIAGo: TIAGo <<SimAgent>>
Figure 12: IBD of GazeboTIAGo: TIAGo <<SimAgent>>
Figure 13: IBD of TIAGo <<PhyAgent>>
Structure evaluation
The final structure evaluation resulted with: IIF=1, MIFRobot=1, MIFFallDetector=1, DGF=1, and DTC=0.67. The result means the structure consists of no <<SimContSubsys>> or <<PhyContSubsys>>, and one <<PhyAgent>> is not mirrored by a DT— SmartHome <<SimAgent>>. The lack of DT for SmartHome <<SimAgent>> results from the requirements (Fig. 10(a)), where Smart Home is not <<SimPhyPartReq>>); thus, as the requirement is not modified, this case is the designer’s informed decision. DGF=1 means the considered SmartHome <<SimAgent>> has <<SimPhyContSubsys>>; therefore, its control subsystem can be shared with a DT, if one will be required in the future. One of the previous design iterations resulted with: IIF=$\frac{5}{7}=0.71$, MIFRobot=1, MIFFallDetector=0, DGF=$\frac{20}{22}=0.91$, DTC=0.67. In this iteration the FallDetector <<SimAgent>> and FallDetector <<PhyAgent>> use embodiment specific control layer (<<SimContSubsys>> and <<PhyContSubsys>>), because FallDetector <<SimAgent>> consists <<SimContSubsys>>, which can not be shared with its DT. To increase MIFFallDetector and DGF the common part of FallDetector <<SimContSubsys>> and FallDetector <<PhyContSubsys>> was extracted. The common part constitutes FallDetector <<SimPhyContSubsys>> in the final design. The FallDetector <<SimPhyContSubsys>> becomes a universal interface between ComplexTaskExecution <<HybAgent>> and the driver layer of FallDetector <<PhyAgent>>. It forced decomposition of FallDetector <<SimAgent>> to FallDetector <<SimPhyContSubsys>>, FallDetector <<SimVirtRec>> and FallDetector <<SimRealRec>>. The latter two simulate Fall Detector sensing. Thanks to the design evaluation, FallDetector <<SimPhyContSubsys>> is used in both embodiments in the final structure, and this part of FallDetector will be tested in simulation because it is common in simulated and physical embodiments of the system.
Configurability definition
Based on the <<OptionalReq>>s in the requirements we define 6 <<SysSetupGpAgents>>s composing of:
  • FallDetector <<PhyAgent>> or FallDetector <<SimAgent>> or no FallDetecor, and
  • Robot <<MirrPhyGpAgents>> or Robot <<MirrSimGpAgents>>.
We show Simulated <<SysSetupGpAgents>> (Fig. 14) and Physical <<SysSetupGpAgents>> (Fig. 15) Internal Block Diagrams (IBDs) to exemplify <<SysSetupGpAgents>> specification. Simulated <<SysSetupGpAgents>> presents two realisations of TIAGo <<SimAgent>>GazeboTIAGo and O3deTIAGo.
Figure 14: Simulated <<SysSetupGpAgents>> with example requirement allocations
Figure 15: Physical <<SysSetupGpAgents>> with example requirement allocations
Implementation
In INCARE the Robot <<PhyAgent>> and Robot <<SimAgent>> are realised with TIAGo robot and its simulations. PAL Robotics mostly implemented these. We mapped the robot hardware and software to our design (as shown in Fig. 11, 12 and 13). RosControl <<SimDriverGpSubsys>> is a software package [^6] http://wiki.ros.org/ros_control implementing various common ROS controllers, and SimTIAGoDrives <<SimHardGpSubsys>> is a package [^7] http://wiki.ros.org/pal_hardware_gazebo implementing Gazebo plugins controlling TIAGo drives. Additionally, we improved the robot’s voice communication by integrating a Large Language Model for user request interpretation [38]. We equipped TIAGo <<PhyAgent>> with additional USB Microphones [57] to improve its audio perception. Complex Task Execution <<HybAgent>> is realised with TaskER framework [55, 36]. The simulated and physical worlds’ synchronisation feature (mirroring humans in the Simulated World) was implemented with HuBeRo framework as WorldSync <<WorldMirrGpAgents>> and tested in a crowded hospital environment (Fig. 16).
Figure 16: A frame from WorldSync <<WorldMirrGpAgents>> testing.
Validation
The INCARE system was deployed and validated in an end-user home, and the videos present its performance in the example tasks— transportation (Fig. 2(a)) [^8] https://vimeo.com/670252925 and fall assistance (Fig. 2(b)) [^9] https://vimeo.com/670246589.
6.2 SPSysML assessment
Based on the open-source SPSysML documentation, a third-party developer designed a system for quantitative evaluation of autonomous vacuum cleaner navigation algorithm [58]. This design activity scored the best possible evaluation factor values. The designer is sure the requirement-based system structuring helps in comprehensive analysis and would use it again as it simplifies CPS design. He commented on the SPSysML: SPSysML promotes wider simulation-based V&V of the system’s change, and evaluation factors help validate and point out the system’s weak parts. However, I am unsure how to implement hybrid agents in ROS, e.g., the same communication topics and structure. Early decomposition is the hardest part of the SPSysML application.
To evaluate our methodology wider, we asked systems engineering and robotics practitioners to assess SPSysML from various points of view. The questions regard individual experience Fig. 17, SPSysML utility Fig. 18, factor improvement guidance Fig. 19, drawbacks, difficulties and our methodology coupling possibilities with standard tools like Zachman Framework [59] and V-model [60]. Finally, the responders assessed our methodology against the qualities of great models Fig. 20. The key questionnaire form and all results are published [^10] https://github.com/RCPRG-ros-pkg/spsysml. The example responses include drawbacks:
  • Evaluation factors might be abused as the importance of components is not taken into account
  • The proposed design evaluation factors are not clearly understandable without a good grasp of SPSysML.
  • With a large number of requirements that have a complex structure, the diagrams may be unreadable. The solution would be a requirements association table.
and advantages:
  • SPSysML certainly gives you the opportunity to evaluate the system architecture and make conscious design decisions.
  • At early stage of system development SPSysML can show some issues regarding system design
  • Joining simulated and physical models provides numerous benefits. SPSysML allows to clearly depict this in system architecture.
  • SPSysML is a systematic framework for joining simulation and physical systems. This process can be controlled with SPSysML.
All responders would apply the requirement-based composition, and eight of nine would apply SPSysML for system specification and design evaluation.
Figure 17: Tasks in which the responders utilise simulation.
Figure 18: How SPSysML would help you in change analysis of a complex CPS?
Figure 19: Clarity of the factor-based rules for integrity improvement.
Figure 20: SPSysML assessment against qualities of great models: 1) Linked to Decision Support, 2) Model Credibility, 3) Clear Scope, 4) V&V With Model, 5) Well-Organized, 6) Analyzable and Traceable, 7) Data Extrapolation, 8) Complete, 9) Internally Consistent, 10) Verifiable, 11) Validation, 12) Model Fidelity, 13) Elegant, 14) Formed for Optimization, 15) Avoid Black Box, 16) Availability of Interfaces, 17) Reusable. The qualities are defined in the questionnaire.
7 Discussion
We state that incorporating SPSysML-based quantitative evaluation of system design indicates significant system structure changes resulting from the proposed quantitative factor-based guidelines applications. An example is FallDetector <<SimAgent>> and FallDetector <<PhyAgent>> integration improvement by a common function isolation as <<SimPhyContSubsys>>. Moreover, the verification shows that 33% of the system’s Physical Agents do not have DTs, and thanks to the design modification, the whole hardware (physical and simulated) is controlled with a common component (DGF=1). This has two main advantages. First, all Digital and Physical Twins pairs share a common control subsystem. Second, if there is a <<PhyAgent>> or <<SimAgent>> without a twin, it can be easily integrated using the embodiment-abstract interface.
We state that the requirement profile for SPSys enables critical analysis and may result in requirement decomposition and explication. An example is the decomposition of the Communication with humans <<SimPhyFunReq>> to TTS and STT <<SimPhyFunReq>> and Dialoge management <<CompFunReq>> during requirement analysis. Thanks to the Dialoge management <<CompFunReq>> separation, the proposed requirement-based procedure for the system composition resulted in the implementation of Talker <<HybAgent>> managing the requirement. As a consequence of the above, the requirement is managed by a component usable in different setups in cooperation with the simulated or physical embodiment. Otherwise, the requirement would be managed separately in simulated and physical embodiments of the robot.
Systems engineering practitioners who reviewed SPSysML confirm the above statements and point out minor difficulties in the SPSysML application.
8 Conclusion
Numerous cyber-physical systems include simulated parts such as Digital Twins (DT), demonstrators, or mockups, which are utilised during their development. We postulate the Simulated-Physical System (SPSys) concept to describe this kind of system. SPSys includes physical, simulated and hybrid parts.
SPSys application is wide; in particular, it can work in a dynamic environment and observe exogenous actions of the inhabitants in the environment. Such a situation is problematic because DT must perfectly mirror its PT. To answer the above needs, we propose a Domain-Specific Language named Simulated-Physical System Modeling Language (SPSysML) that defines SPSys taxonomy and the relationships between the types of SPSys parts.
Integrity and reusability in reliable software development is crucial. Therefore, based on SPSysML terms, we define design evaluation factors that enable quantitative analysis and optimisation targeted to software re-use maximisation between DT and PT in different system setups. We analyze the evaluation factors and show features of the systems that score edge case values of the factors.
Finally, we validate SPSysML, including the evaluation factors and proposed requirement-based system composition in a complex robot system development. We demonstrate step-by-step results of SPSysML application. The validation shows that the requirement profile for SPSys enables critical analysis and may result in requirement decomposition and explication. The validation also demonstrates the positive influence of the SPSysML-based quantitative evaluation on component reusability and integrity between simulated and physical parts.
The presented validation confirms the features distinguishing our method from the others listed in Table 1, e.g.:
  • Quantitative evaluation of integrity between the system’s simulated and physical embodiments,
  • Modeling simulation, physical and hybrid system execution and testing setups,
  • SysML-based specification of systems composed of physical and simulated parts and Digital Twins.
  • Enabling Digital Twins to observe exogenous actions in the simulated world by <<WorldMirrGpAgents>> implementation.
  • Easing requirements tracing using requirement-based system structuring.
Based on the conducted validation, we observe the following limitations of our work:
  • Requirement-based structuring is a roadmap for SPSys’s design, increasing structural connections between system components and requirements. However, it may not be optimal for any system; therefore, we suggest further iterating the design stage with other system design methods, like those based on the Design Structure Matrix [61].
  • The requirement model defines only types used in the requirement-based structuring method; however, the project team should extend the requirement model if other requirement types are necessary.
  • The integrity evaluation accuracy depends on the system structure granularity, which is the resolution of the evaluation. Therefore, the more detailed the system design is, the more accurate the evaluation.
  • The drawbacks pointed out in the questionnaire. Evaluation factors are objective and abstract from the subjective importance of particular components. However, such aspects can be applied by scaling the proposed factors by importance, priority, etc.
In the future, we plan to automate SPSys structure evaluation with known universal tools like Matlab and Enterprise Architect. Furthermore, this work revealed the need for a CPS structure optimisation indicators taxonomy. SPSysML and future work lead to optimal structure development automation for complex robot systems like dual-arm impedance controlled mobile manipulatiors [10].
Acknowledgment
The research was funded by the Centre for Priority Research Area Artificial Intelligence and Robotics of Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. The work takes the robot platform and its application from the INCARE project. The authors also acknowledge TALBOT, from the European Union’s Horizon 2020 research and innovation programme under Marie Skłodowska-Curie grant agreement No. 801342 (Tecniospring INDUSTRY) and the Government of Catalonia’s Agency for Business Competitiveness (ACCIÓ); and SHAPES, from the European Horizon 2020 research and innovation programme under grant agreement No 857159. The Spanish grant PID2021-125535NB-I00 has also supported the work.